Use Boyle's, Charles's, Gay-Lussac's Law to calculate the missing value of the following:

v1= 250 ml, t1=?, v2=400 ml, t2=298 Kelvin
1
Expert's answer
2018-01-22T05:57:35-0500
T1/V1 = T2/V2
T1 = (V1·T2)/V2
T1 = (250ml·298K)/400ml = 186.25 K